OpenStudy (anonymous):

Determine Slope-Int Form of a line containing point (.5,1.5) and slope -1, y=mx+b, right?

OpenStudy (anonymous):

That's the form, yes. The point is to think "Oh, at x = 0.5, y = 1.5 and m = -1." Put that into the form y = mx + b and solve for b.

OpenStudy (anonymous):

oh I have to solve for b? I thought I was suppose to do this y - 1.5 = -1x - .5

OpenStudy (anonymous):

doesn't that work too?

OpenStudy (anonymous):

You have m. You have a point on the graph. There are several lines which can have those properties. What determines a line is its slope and y-intercept. To find the y-intercept (b), you solve for it in the equation y = mx + b. You have y, x, and m. You do that by saying "Oh. b = y - mx = 1.5 - m*0.5". You see where this is going ... ? I *think* what you're doing is point-slope form, so find an answer using that. Compare your answer with what you get when you use my method. See if that works! :)

OpenStudy (anonymous):

Oh I see, same answer to different forms ,right?

OpenStudy (anonymous):

*two

OpenStudy (anonymous):

Yep. :)